# Gated Linear Attention Transformers with Hardware-Efficient Training

### 12 Dec 2023 v2

@@ 1 Introduction
In this work, we derive an efficient training strategy for linear attention Transformers with data-dependent forget gates. We first derive a parallel form of our gated linear attention (GLA) layer, similar to the parallel form of softmax attention. With this form, the RNN outputs can be computed in parallel without materializing the 2D hidden states for all tokens. However, this parallel form by itself is not amenable to tensor core accelerations due to some non-standard matrix multiplication operations that are needed for numerical stability. We further develop a chunk-wise parallel form of the GLA layer, which strikes a balance between the recurrent form and the parallel form. This approach employs chunk-level recurrence to sync information between chunks, while using the parallel form for intra-chunk computations. With this form we can leverage standard half-precision matrix matmuls for most computations, thus enabling efficent training.
On moderate-scale language modeling experiments, where we train models with 350M/1.3B parameters on 15B/100B tokens respectively, we find that the GLA Transformer performs on par with a (1) strong LLaMA-architecture Transformer baseline which make use of recent recipes (${}^{++}$; Touvron et al., 2023) as well as (2) Mamba (Gu & Dao, 2023), a recent state-space models that uses data-dependent state transition mechanisms. For training efficiency, our Triton-based implementation (Tillet et al., 2019) performs comparably to CUDA-optimized FlashAttention-2 (Dao, 2023) under the regular 2048 training length setting, while outperforming FlashAttention-2 when training on longer sequences beyond 4096.
@@ 2.1 Linear Attention Transformers
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+\phi({\mathbf{K}}_{i})^{\cramped{{}^{\mathsf{T}}}}{\mathbf{V}}_{i},\qquad{\bm{z}}_{t}={\bm{z}}_{t-1}+\phi({\mathbf{K}}_{i})^{\cramped{{}^{\mathsf{T}}}},\qquad{\mathbf{O}}_{t}=\frac{\phi({\mathbf{Q}}_{t}){\mathbf{S}}_{t}}{\phi({\mathbf{Q}}_{t}){\bm{z}}_{t}}.$
@@ 3.1 Recurrent Form
${\mathbf{G}}_{t}$ $=\bm{\alpha}_{t}^{\cramped{{}^{\mathsf{T}}}}\bm{\beta}_{t}\in\mathbb{R}^{d_{k}\times d_{v}},\qquad\bm{\alpha}=\sigma(({\mathbf{X}}{\bm{W}}_{\alpha}+{\bm{b}}_{\alpha}))/\tau\in\mathbb{R}^{L\times d_{k}},\quad\bm{\beta}=\sigma(({\mathbf{X}}{\bm{W}}_{\beta}+{\bm{b}}_{\beta}))/\tau\in\mathbb{R}^{L\times d_{v}},$
where $\sigma$ is the Sigmoid function and $\tau\in\mathbb{R}$ is a temperature term.888 Incorporating $\tau>1$ is to mitigate against the sigmoid outputs quickly saturating. In our experiments, we set $\tau=16$. This outer-product-based parameterization of ${\mathbf{G}}_{t}$ is parameter-efficient [^9] Since it only requires $O(d\times d_{k}+d\times d_{v})$ parameters, while a full parameterization would need $O(d\times d_{k}\times d_{v})$. and it moreover has the same outer-product structure as the RNN input ${\mathbf{K}}_{t}^{\cramped{{}^{\mathsf{T}}}}{\mathbf{V}}_{t}$ which, as we will shortly see, allow for equivalent parallel forms. The output of a GLA layer for time step $t$ is
@@ 3.4 GLA-Transformer
${\mathbf{Y}}^{(l)}=\text{GLA}(\text{LN}({\mathbf{X}}^{(l)}))+{\mathbf{X}}^{(l)},\qquad{\mathbf{X}}^{(l+1)}=\text{SwiGLU}({\mathbf{Y}}^{(l)}),$
@@
${\bm{\alpha}}=\sigma(({\mathbf{X}}{\bm{W}}^{1}_{\alpha}{\bm{W}}^{2}_{\alpha}+{\bm{b}}_{\beta}))/\tau\in\mathbb{R}^{L\times d_{v}},$

### 24 Dec 2023 v3

@@ 1 Introduction
In this work, we derive an efficient training strategy for linear attention Transformers with data-dependent forget gates. We first derive a parallel form of our gated linear attention (GLA) layer, similar to the parallel form of softmax attention. With this form, the RNN outputs can be computed in parallel without materializing the 2D hidden states for all tokens. However, this parallel form by itself is not amenable to tensor core accelerations due to some non-standard matrix multiplication operations that are needed for numerical stability. We further develop a chunk-wise parallel form of the GLA layer, which strikes a balance between the recurrent form and the parallel form. This approach employs a chunk-level recurrence to sync information between chunks, while using the parallel form for intra-chunk computations. With this form, we can leverage standard half-precision matrix matmuls for most computations, thus enabling efficient training.
On moderate-scale language modeling experiments, where we train models with 340M/1.3B parameters on 15B/100B tokens respectively, we find that the GLA Transformer performs on par with a (1) strong LLaMA-architecture Transformer baseline which makes use of recent recipes (${}^{++}$; Touvron et al., 2023) as well as (2) Mamba (Gu & Dao, 2023), a recent state-space models that uses data-dependent state transition mechanisms. For training efficiency, our Triton-based implementation (Tillet et al., 2019) performs comparably to CUDA-optimized FlashAttention-2 (Dao, 2023) under the regular 2048 training length setting, while outperforming FlashAttention-2 when training on longer sequences beyond 4096.
@@ 2.1 Linear Attention Transformers
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+\phi({\mathbf{K}}_{t})^{\cramped{{}^{\mathsf{T}}}}{\mathbf{V}}_{t},\qquad{\bm{z}}_{t}={\bm{z}}_{t-1}+\phi({\mathbf{K}}_{t})^{\cramped{{}^{\mathsf{T}}}},\qquad{\mathbf{O}}_{t}=\frac{\phi({\mathbf{Q}}_{t}){\mathbf{S}}_{t}}{\phi({\mathbf{Q}}_{t}){\bm{z}}_{t}}.$
@@ 3.1 Recurrent Form
${\mathbf{G}}_{t}$ $=\bm{\alpha}_{t}^{\cramped{{}^{\mathsf{T}}}}\bm{\beta}_{t}\in\mathbb{R}^{d_{k}\times d_{v}},\qquad\bm{\alpha}=\sigma({\mathbf{X}}{\bm{W}}_{\alpha}+{\bm{b}}_{\alpha})/\tau\in\mathbb{R}^{L\times d_{k}},\quad\bm{\beta}=\sigma({\mathbf{X}}{\bm{W}}_{\beta}+{\bm{b}}_{\beta})/\tau\in\mathbb{R}^{L\times d_{v}},$
where $\sigma$ is the Sigmoid function and $\tau\in\mathbb{R}$ is a temperature term.888 Incorporating $\tau>1$ is to mitigate against the sigmoid outputs quickly saturating. In our experiments, we apply the temperature in log-space: $\log{\bm{\alpha}}=\big{(}\log\sigma({\mathbf{X}}{\mathbf{W}}_{\alpha}+{\bm{b}}_{\alpha}))/16,\,\log{\bm{\beta}}=\big{(}\log\sigma({\mathbf{X}}{\bm{W}}_{\beta}+{\bm{b}}_{\beta}))/16$. This outer-product-based parameterization of ${\mathbf{G}}_{t}$ is parameter-efficient [^9] Since it only requires $O(d\times d_{k}+d\times d_{v})$ parameters, while a full parameterization would need $O(d\times d_{k}\times d_{v})$. and it moreover has the same outer-product structure as the RNN input ${\mathbf{K}}_{t}^{\cramped{{}^{\mathsf{T}}}}{\mathbf{V}}_{t}$ which, as we will shortly see, allow for equivalent parallel forms. The output of a GLA layer for time step $t$ is
@@ 3.4 GLA-Transformer
${\mathbf{Y}}^{(l)}=\text{GLA}(\text{LN}({\mathbf{X}}^{(l)}))+{\mathbf{X}}^{(l)},\qquad{\mathbf{X}}^{(l+1)}=\text{SwiGLU}(\text{LN}({\mathbf{Y}}^{(l)})),$
@@
${\bm{\alpha}}=\sigma({\mathbf{X}}{\bm{W}}^{1}_{\alpha}{\bm{W}}^{2}_{\alpha}+{\bm{b}}_{\alpha})/\tau\in\mathbb{R}^{L\times d_{k}},$